A particle starts moving on a circular direction with angular speed 5 rads−1. The radius of the circle is 4 m and the angular speed at time t is given by ω =15−3t.

What is the linear speed of the particle 6 seconds after moving, and the resultant acceleration of the particle?


1
Expert's answer
2021-06-10T03:02:13-0400

Ans:-

The relationship between angular velocity ω and linear velocity v was also defined in Rotation Angle and Angular Velocity as

"v=r\times \omega," "v=r\times \omega"

We have "\omega =15-3t" and "r=4"

So, "v=4 \u00d7 (15-3t)"

linear speed of the particle 6 seconds after moving is

"t=6 \Rightarrow v=4\times (15-3\times6)=-12 \ \ m/s" "-ve" sign means direction of speed is in opposite direction of motion.

"\alpha=\dfrac{d\omega}{dt}=15"


So, tangential acceleration

"a_t=\alpha\times r=15\times4=60"


Centripetal acceleration will be

"a_c=\omega^2\times r=5^2\times4=100"


Then resultant acceleration will be


"a=\sqrt{60^2+100^2}=116.61 m/s^2"


Hence the resultant acceleration of the particle is "116.61 \ \ m/s^2"